Use a random number generator to simulate the roll of a fair die 100 times. Let the number face up on the die represent the variable X

A. Build a relative frequency table of the outcomes of the variable X.

X

Freq

Rel. Freq

B. Use the relative frequency distribution from part c to estimate the probability of an even number face up, then find the actual probability using the probability distribution and comment on the difference in values.
Answer #1
throws <- sample(1:6, 100, replace=TRUE, prob=c(1,1,1,1,1,1)/6 )

table(throws)
throws
1 2 3 4 5 6
13 21 14 19 15 18

x f rel.freq.
1 13 0.13
2 21 0.21
3 14 0.14
4 19 0.19
5 15 0.15
6 18 0.18

probability of even number using relative frequency distribution = 0.21+0.19+0.18=0.58

actual probability = 0.5